# What Are They Filtering Out? A Survey of Filtering Strategies for Harm Reduction in Pretraining Datasets

### v1

@@ Abstract
Data filtering strategies are a crucial component to develop safe Large Language Models (LLM), since they support the removal of harmful contents from pretraining datasets. There is a lack of research on the actual impact of these strategies on vulnerable groups to discrimination, though, and their effectiveness has not been yet systematically addressed. In this paper we present a benchmark study of data filtering strategies for harm reduction aimed at providing a systematic overview on these approaches. We survey $55$ technical reports of English LMs and LLMs to identify the existing filtering strategies in literature and implement an experimental setting to test their impact against vulnerable groups. Our results show that the positive impact that strategies have in reducing harmful contents from documents has the side effect of increasing the underrepresentation of vulnerable groups to discrimination in datasets.
@@
What Are They Filtering Out? A Survey of Filtering Strategies for Harm Reduction in Pretraining Datasets
Marco Antonio Stranisci Università degli Studi di Torino aequa-tech marcoantonio.stranisci@unito.it                        Christian Hardmeier IT University of Copenhagen chrha@itu.dk
@@ 2 Related work
Research on data filtering strategies for harm reduction has been propelled by the research on biases in LLMs Bender et al. (2021) and in pretraining datasets Jo and Gebru (2020). However, critical studies on strategies focused on specific datasets rather than the impact of strategies themselves. Luccioni and Viviano (2021) analyzed the presence of Hate Speech in the Common Crawl dataset. Dodge et al. (2021) documented the C4 corpus Raffel et al. (2020), showing that its filtering correlates with a reduction of terms defining vulnerable groups to discrimination in datasets. Longpre et al. (2024) tested the correlation between removing toxic contents from pretraining datasets and LLMs performance in toxicity classification, showing that filtering has a negative impact on this task. The survey of Albalak et al. (2024) on data selection for LLMs provide a marginal and non-systematic analysis of strategies. Finally, the work of Lucy et al. (2024) on language filtering strategies provide insightful results on the cultural biases embedded in language filters that can determine the exclusion of certain categories of people from pretraining datasets.
3 Surveying Data Filtering Strategies
In this section we present our systematic survey of data filtering strategies for harm reduction in pretraining data as they are presented in $55$ LMs technical reports. We first introduce our methodology for the selection of relevant documents and their analysis. Then, we present a taxonomy of eight existing data filtering strategies that emerges from the analysis. Finally, we identify some trends emerging from a general overview of reports in a diachronic perspective.
3.1 Survey Methodology
@@
document seeding GPT-3 Brown et al. (2020), DSIR Xie and Agrawal (2023), Mamba Gu and Dao (2023), MPT-7B Research (2023), MAmmoTH2 Mickus et al. (2024)
@@
safety policy Gemini Gemini et al. (2023), EXAONE 3.5 Research et al. (2024), INCITE-RedPajama Weber et al. (2024), LLama-3 Dubey (2024)
@@
For each report we checked three types of content: the description of the pretraining dataset, the presentation of data filtering strategies, and the treatment of harmful or biased contents in datasets [^2] In the context of this research, we adopt the term harmful for all the phenomena that might have a negative impact against vulnerable groups to discrimination. In this sense we refer to survey of Blodgett et al. (2020), who distinguish between representational harms, which consist in all the negative representation of groups, and allocative harms, which includes all the forms of underrepresentation. For each paper we search the following keywords: data, filter, quality, toxic, bias, hate, stereotype. This stage led to the retrieval of $13$ additional papers referenced in the sections that cover the description of filtering strategies. Among them $4$, describe the creation of a pretraining dataset and $1$ a methodology for data sampling. At the end of this process we obtained a pool of $55$ papers for our survey with information about the type of data filtering strategies that have been implemented to remove harmful contents and their availability.
@@ 3.3 Trends and Limitations in Data Filtering
In this section we discuss some general trends and limitations that emerge from our survey.
A first consideration is about the the lack of replicability of filtering strategies presented in technical reports. This trend is shown in Table 1 where models are highlighted in red if they did not disclose their data filtering pipeline or the snapshot of datasets used for training, in orange if they partially did it, in green if they fully disclose filtering strategies and data. As can be observed, most of the existing LMs are not released with the actual scripts for the replication of implemented filtering methods. Specifically, we recognize a first generation of LMs (e.g.: BERT, RoBERTa, etc.) that have been trained without considering the issue of harmful contents in pretraining datasets. The pivotal work of Bender et al. (2021) led to a second generation of models that implemented specific strategies and open-sourced their approaches and results for further investigation. This is the case of the C4 corpus Raffel et al. (2024), which has been released both in a filtered and unfiltered version. Besides notable exceptions Soldaini et al. (2024); Weber et al. (2024), the current tendency is not to disclose scripts and results of the data filtering strategies implemented for pretraining.
A similar trend towards the reduced openness of filtering strategies is observable about the documentation debt of implemented filtering strategies. If the introduction of documentation templates like Datasheets Gebru et al. (2021) and Model Cards Mitchell et al. (2019) led to a first wave of fully documented self-assessment reports Chowdhery et al. (2023), the more recent technical reports formally rely on these documentation templates but substantially provide very generic descriptions about their data filtering approach. For instance, the description provided in the Gemma technical report is limited to the following paragraph: “We filter the pre-training dataset to reduce the risk of unwanted or unsafe utterances […]. This includes both heuristics and model-based classifiers to remove harmful or low-quality content” Team et al. (2024a).
@@ 4.1 Experimental Setting
Table 2: The average number of named entities in the 5 samples gathered from Common Crawl and the percentage of sentences removed by applying $7$ filtering strategies. Named entities are broken down by groups resulting from the intersection of gender and origin: Western men (w.m.), Post-colonial men (p-c.m.), Western women (w.-w.), and Post-colonial women (p-c.-w.)
@@
The first step of our experimental setting was the creation of a knowledge base that enables the categorization of named entities on the basis of their gender and origin. In order to do so we developed the People Dataset, a corpus of $10.8$ million of entities of the type person with information about their country of birth, citizenship, gender, ethnic minority, and occupation. The dataset is derived from Wikidata Erxleben et al. (2014), a collaborative Knowledge Graph (KG) maintained by the Wikimedia ecosystem, and processed on the basis of previous literature on the topic Stranisci et al. (2023): we inferred people’s countries of birth from Wikidata property ‘place of birth’ (P19) properties, mapped all the properties of the type ‘ethnic group’ (P172) with a curated list of Post-colonial minorities in Western countries (e.g., African-Americans). Finally, we integrated the knowledge base with additional information from CaLiGraph Heist and Paulheim (2019), a KG derived from Wikipedia categories.
@@
We gathered $5$ samples of documents from the Common Crawl (CC) snapshot released in August 2024666https://data.commoncrawl.org/crawl-data/CC-MAIN-2024-33/index.html. Each sample is composed of $20$ Web ARChive (WARC) files that have been processed according to the following criteria: i. we kept only documents classified as written in English, implementing the langdetect Python library [^7] https://pypi.org/project/langdetect/; ii. we removed all the documents with an average number of sentences below x and an average number of words per sentence below y. The average number of documents in each sample is $102,820$ (std: $443$).
@@
The next step of our approach was the linking of named entities with the People Dataset. In order to adopt a computationally-efficient Entity Linking (EL) approach to implement on a large set of data, we designed an EL pipeline that combines the adoption of neural methods for the detection of named entities and existing heuristics to link them to the dataset Manghi (2023). The pipeline is organized in four steps. i. We used the largest SpaCy model [^8] https://spacy.io/,en_core_web_sm to detect all the named entities of the type PERSON in CC samples. ii. We queried Wikipedia APIs [^9] https://www.mediawiki.org/wiki/API:Search with all the retrieved named entities to obtain titles of Wikipedia pages and filtered out all the named entities that were too dissimilar from Wikipedia titles, through the adoption of a set of heuristics already validated by Manghi (2023). iii. We retrieved all the Wikidata IDs corresponding to the Wikipedia titles. iv. We kept only entities with a Wikidata ID that is present in our People Dataset. As a result (Table 2), we identified an average of $128,025$ entities from each sample: $85,276$ Western men, $15,272$ Post-colonial men, $22,185$ Western women, $5,292$ Post-colonial women. In order to assess whether there is high variation between samples we conducted an ANOVA test St et al. (1989) on all the possible pairs of entities distributions broken-down by group. In all cases we obtained a $p-value$ of $0.99$, showing that there is no evidence of variation between different CC samples.
@@
Having created our baseline, we implemented seven filtering strategies belonging to three different approaches that can be replicated in an experimental setting (Section 3.2): rule-based, toxicity classifier, and document similarity. These strategies were applied to all the sentences or documents that mentioned at least one of the retrieved named entities.
@@ 5 Discussion
In contrast with this evidence about the complex nature of data filtering for harm reduction, the general tendency that emerges from our survey (RQ 1) is disengagement with this issue. After a wave of efforts in this field connected with the critical work of Bender et al. (2021), the interest in implementing robust pretraining data filtering dramatically reduced in favor of post-training measures for harm reduction. In most cases, the actual implementation of strategies is close-sourced, hindering the replication of procedures that are adopted for processing datasets. The presence of notable exceptions such as Dolma Soldaini et al. (2024) and the RefinedWeb Dataset Penedo et al. (2023) represents a significant countertendency. However, there is still a lack of filtering strategies that are effective in balancing the need to remove harmful contents while preserving the representativeness of vulnerable groups in pretraining datasets.
@@ 6 Conclusion
In this paper we presented a first systematic analysis of data filtering strategies for harm reduction in datasets. The survey of $55$ technical reports describing LLMs enabled us to draw a taxonomy of filtering strategies and to identify the open issues that prevent the implementation of effective, reliable, and replicable methods for the reduction of harmful contents. Additionally, we benchmarked seven existing strategies on CC samples of documents in order to assess their impact against groups characterized by the intersection of gender and origin. The benchmark showed a systematic negative effect of strategies against women and a high variety in the types of harmful contents that specific strategies filter out.
Future work will focus on providing a more effective pipeline for the assessment of data filtering strategies with two aims in mind: i. improving the coverage of our pipeline by including additional information regarding groups that does not rely only on named entities; ii. adopting a participatory approach to the evaluation of filtering strategies that engages associations and communities of people who are active against discrimination. Finally, we will also systematically explore the correlation between the adoption of specific data filtering strategies and linguistic behaviors emerging from models trained on filtered datasets.

### v2

# What Are They Filtering Out? An Experimental Benchmark of Filtering Strategies for Harm Reduction in Pretraining Datasets

@@ Abstract
Data filtering strategies are a crucial component to develop safe Large Language Models (LLM), since they support the removal of harmful contents from pretraining datasets. There is a lack of research on the actual impact of these strategies on vulnerable groups to discrimination, though, and their effectiveness has not been yet systematically addressed. In this paper we present a benchmark study of data filtering strategies for harm reduction aimed at providing a systematic evaluation on these approaches. We provide an overview $55$ technical reports of English LMs and LLMs to identify the existing filtering strategies in literature and implement an experimental setting to test their impact against vulnerable groups. Our results show that the positive impact that strategies have in reducing harmful contents from documents has the side effect of increasing the underrepresentation of vulnerable groups to discrimination in datasets.
@@
What Are They Filtering Out? An Experimental Benchmark of Filtering Strategies for Harm Reduction in Pretraining Datasets
Marco Antonio Stranisci University of Turin aequa-tech marcoantonio.stranisci@unito.it          Christian Hardmeier IT University of Copenhagen chrha@itu.dk
@@ 2 Related work
Research on data filtering strategies for harm reduction has been propelled by the research on biases in LLMs Bender et al. (2021) and in pretraining datasets Jo and Gebru (2020). However, critical studies on strategies focused on specific datasets rather than the impact of strategies themselves. Luccioni and Viviano (2021) analyzed the presence of Hate Speech in the Common Crawl dataset. Dodge et al. (2021) documented the C4 corpus Raffel et al. (2020), showing that its filtering process correlates with a reduction of terms defining vulnerable groups to discrimination in datasets. Longpre et al. (2024) tested the correlation between removing toxic contents from pretraining datasets and LLMs performance in toxicity classification, showing that filtering has a negative impact on this task. The survey of Albalak et al. (2024) on data selection for LLMs provide a marginal and non-systematic taxonomy of strategies. Finally, the work of Lucy et al. (2024) on language filtering strategies provide insightful results on the cultural biases embedded in language filters that can determine the exclusion of categories of people who speak non-standard English varieties.
Our work fills the existing gap providing a first systematic analysis of the impact of filtering strategies for harm reduction in datasets.
3 An Overview of Existing Data Filtering Strategies
In this section we present our systematic overview of data filtering strategies for harm reduction in pretraining data as they are presented in $55$ LMs technical reports. We first introduce our methodology for the selection of relevant documents and their analysis. Then, we present a taxonomy of eight existing data filtering strategies that emerges from the analysis. Finally, we identify some trends emerging from a general overview of reports in a diachronic perspective.
3.1 Overview Methodology
@@
document seeding GPT-3 Brown et al. (2020), DSIR Xie et al. (2023), Mamba Gu and Dao (2023), MPT-7B Research (2023), MAmmoTH2 Mickus et al. (2024)
@@
safety policy Gemini Gemini et al. (2023), EXAONE 3.5 Research et al. (2024), INCITE-RedPajama Weber et al. (2024), LLama-3 Dubey et al. (2024)
@@
For each report we checked three types of content: the description of the pretraining dataset, the presentation of data filtering strategies, and the treatment of harmful or biased contents in datasets. In the context of this research, we adopt the term harmful for all the phenomena that might have a negative impact against vulnerable groups to discrimination. In this sense we refer to survey of Blodgett et al. (2020), who distinguish between representational harms, which consist in all the negative representation of groups, and allocative harms, which includes all the forms of underrepresentation.
For each paper we search the following keywords: data, filter, quality, toxic, bias, hate, stereotype. This stage led to the retrieval of $13$ additional papers referenced in the sections that cover the description of filtering strategies. Among them, $4$ describe the creation of a pretraining dataset and $1$ a methodology for data sampling. At the end of this process we obtained a pool of $55$ papers for our survey with information about the type of data filtering strategies that have been implemented to remove harmful contents and their availability.
@@ 3.3 Trends and Limitations in Data Filtering
In this section we discuss some general trends and limitations that emerge from our overview.
A first consideration is about the the lack of replicability of filtering strategies presented in technical reports. This trend is shown in Table 1 where models are highlighted in red if they did not disclose their data filtering pipeline or the snapshot of datasets used for training, in orange if they partially did it, in green if they fully disclose filtering strategies and data. As can be observed, most of the existing LMs are not released with the actual scripts for the replication of implemented filtering methods. Specifically, we recognize a first generation of LMs (e.g.: BERT, RoBERTa, etc.) that have been trained without considering the issue of harmful contents in pretraining datasets. The pivotal work of Bender et al. (2021) led to a second generation of models that implemented specific strategies and open-sourced their approaches and results for further investigation. This is the case of the C4 corpus Raffel et al. (2024), which has been released both in a filtered and unfiltered version. Besides notable exceptions Soldaini et al. (2024); Weber et al. (2024), the current tendency is to not disclose scripts and results of the data filtering strategies implemented for pretraining.
A similar trend towards the reduced openness of filtering strategies is observable about the documentation debt Bandy and Vincent (2021) of implemented filtering strategies. If the introduction of documentation templates like Datasheets Gebru et al. (2021) and Model Cards Mitchell et al. (2019) led to a first wave of fully documented self-assessment reports Chowdhery et al. (2023), the more recent technical reports formally rely on these documentation templates but substantially provide very generic descriptions about their data filtering approach. For instance, the description provided in the Gemma technical report is limited to the following paragraph: “We filter the pre-training dataset to reduce the risk of unwanted or unsafe utterances […]. This includes both heuristics and model-based classifiers to remove harmful or low-quality content” Team et al. (2024a).
@@ 4.1 Experimental Setting
Table 2: The average number of named entities in the 5 samples gathered from Common Crawl and the percentage of sentences removed by applying 77 filtering strategies. Named entities are broken down by groups resulting from the intersection of gender and origin: Western men (w.m.), Post-colonial men (p-c.m.), Western women (w.-w.), and Post-colonial women (p-c.-w.)
@@
The first step of our experimental setting was the creation of a knowledge base that enables the categorization of named entities on the basis of their gender and origin. In order to do so we developed the People Dataset, a corpus of $10.8$ million of entities of the type person with information about their country of birth, citizenship, gender, ethnic minority, and occupation. The dataset is derived from Wikidata Erxleben et al. (2014), a collaborative Knowledge Graph (KG) maintained by the Wikimedia ecosystem, and processed on the basis of previous literature on the topic Stranisci et al. (2023b): we inferred people’s countries of birth from Wikidata property ‘place of birth’ (P19) properties, mapped all the properties of the type ‘ethnic group’ (P172) with a curated list of Post-colonial minorities in Western countries (e.g., African-Americans). Finally, we integrated the knowledge base with additional information from CaLiGraph Heist and Paulheim (2019), a KG derived from Wikipedia categories.
@@
We gathered $5$ samples of documents from the Common Crawl (CC) snapshot released in August 2024555https://data.commoncrawl.org/crawl-data/CC-MAIN-2024-33/index.html. Each sample is composed of $20$ Web ARChive (WARC) files that have been processed according to the following criteria: i. we kept only documents classified as written in English, implementing the langdetect Python library [^6] https://pypi.org/project/langdetect/; ii. we removed all the documents with an average number of sentences below $5$ and an average number of words per sentence below $5$. The average number of documents in each sample is $102,820$ (std: $443$).
@@
The next step of our approach was the linking of named entities with the People Dataset. In order to adopt a computationally-efficient Entity Linking (EL) approach to implement on a large set of data, we designed an EL pipeline that combines the adoption of neural methods for the detection of named entities and existing heuristics to link them to the dataset Stranisci et al. (2023a). The pipeline was organized in four steps.
  1. 1.
    We used the largest SpaCy model [^7] https://spacy.io/,en_core_web_sm to detect all the named entities of the type PERSON in CC samples.
  2. 2.
    We queried Wikipedia APIs [^8] https://www.mediawiki.org/wiki/API:Search with all the retrieved named entities to obtain titles of Wikipedia pages and filtered out all the named entities that were too dissimilar from Wikipedia titles, through the adoption of a set of heuristics already validated by Manghi (2023).
  3. 3.
    We retrieved all the Wikidata IDs corresponding to the Wikipedia titles.
  4. 4.
    We kept only entities with a Wikidata ID that is present in our People Dataset.
As a result (Table 2), we identified an average of $128,025$ entities from each sample: $85,276$ Western men, $15,272$ Post-colonial men, $22,185$ Western women, $5,292$ Post-colonial women. In order to assess whether there is high variation between samples we conducted an ANOVA test St et al. (1989) on all the possible pairs of entities distributions broken-down by group. In all cases we obtained a pvaluep-value of $0.99$, showing that there is no evidence of variation between different CC samples.
@@
Having created our baseline, we implemented seven filtering strategies belonging to three different approaches outlined in the taxonomy 1, which can be replicated in an experimental setting: rule-based, toxicity classifier, and document similarity. These strategies were applied to all the sentences or documents that mentioned at least one of the retrieved named entities.
@@ 5 Discussion
In contrast with this evidence about the complex nature of data filtering for harm reduction, the general tendency that emerges from our survey (RQ 1) is disengagement with this issue. After a wave of efforts in this field connected with the critical work of Bender et al. (2021), the interest in implementing robust pretraining data filtering dramatically reduced in favor of post-training measures for harm reduction. In most cases, the actual implementation of strategies is close-sourced, hindering the replication of procedures that are adopted for processing datasets. The presence of notable exceptions such as Dolma Soldaini et al. (2024) and the RefinedWeb Dataset Penedo et al. (2023) represents a significant counter-tendency. However, there is still a lack of filtering strategies that are effective in balancing the need to remove harmful contents while preserving the representativeness of vulnerable groups in pretraining datasets.
@@ 6 Conclusion
In this paper we presented a first systematic analysis of data filtering strategies for harm reduction in datasets. The overview of $55$ technical reports describing LLMs enabled us to draw a taxonomy of filtering strategies and to identify the open issues that prevent the implementation of effective, reliable, and replicable methods for the reduction of harmful contents. Additionally, we experimentally evaluated seven existing strategies on CC samples of documents in order to assess their impact against groups characterized by the intersection of gender and origin. The evaluation showed a systematic negative effect of strategies against women and a high variety in the types of harmful contents that specific strategies filter out.
Future work will focus on providing a more effective pipeline for the assessment of data filtering strategies with three aims in mind: i. improving the coverage of our pipeline by including additional information in the People Dataset. We plan to include information from additional knowledge bases and add information about vulnerable groups that goes beyond named entities (e.g., mentions of communities, ethnic groups, and demonyms) ; ii. adopting a participatory approach to the evaluation of filtering strategies that engages associations and communities of people who are active against discrimination. This approach will enable to explore the data filtering process from the perspectives of people who might actually harmed by the implementation of strategies; iii we will systematically explore the correlation between the adoption of specific data filtering strategies and linguistic behaviors emerging from models trained on filtered datasets.